OpenStudy (anonymous):

A linear function has a slope of -7/9 and a y-intercept of 3. How does this function compare to the linear function that is represented by the equation y+11=-7/9(x-8) It has the same slope and the same y-intercept. It has the same slope and a different y-intercept. It has the same y-intercept and a different slope. It has a different slope and a different y-intercept.

OpenStudy (paxpolaris):

\[y+11=-\frac79(x-8) \]^^^ what is the slope of the line?

OpenStudy (anonymous):

The first linear equation is: y=-7/9x +3 the second one is:y+11 = -7/9(x-8) or y=-11-7/9(x-8) or y=-7/8x - 51/11 Therefore, they have the same slope but different y-intercept

OpenStudy (anonymous):

Sorry! I meant: y=-7/9x - 51/11 instead of : y=-7/8x -51/11